OpenStudy (anonymous):

What is the equation of the line that passes through (–2, –3) and is perpendicular to 2x – 3y = 6?

OpenStudy (luigi0210):

Perpendicular means the slope is \(-\frac{1}{m}\). SO put the equation in slope intercept-form first. Then use the point-slope formula to get the equation \(y-y_{1}=m(x-x_{1})\)

OpenStudy (anonymous):

y=mx=b

OpenStudy (luigi0210):

Yes, turn \(2x-3y=6\) into \(y=mx+b\)

OpenStudy (anonymous):

y=-2/3x+2/3

OpenStudy (luigi0210):

Actually it's just \( y=-\frac{2}{3}x-2\) but the constant doesn't really matter at this moment, just the slope. So take that slope and find the negative reciprocal of it.

OpenStudy (anonymous):

3/2

OpenStudy (luigi0210):

Correct! Now, we have the slope \(m=\frac{3}{2}\) and the point (\(-2, -3)\). Use the point-slope formula to find the new equation.

OpenStudy (anonymous):

y=3/2x

OpenStudy (luigi0210):

Let's see.. \[\LARGE y+3=\frac{3}{2}(x+2)\] \[\LARGE y+3=\frac{3}{2}x+3\] Yup, right!

OpenStudy (anonymous):

thax i got another

OpenStudy (anonymous):

Write the equation of the line that passes through (–4, 3) and is perpendicular to the line 2003-10-05-00-00_files/i0220000.jpg.

OpenStudy (luigi0210):

Could you try the attachment again?

OpenStudy (luigi0210):

Could you type it? It's not showing :l

OpenStudy (anonymous):

y=7/5x-3/5

OpenStudy (luigi0210):

Same thing, take the slope and get the negative rec. and use point slope.